A quantity a varies directly as x and another quantity b varies inversely as x. when x = 2, then sum of a and b is 7, and when x = 3, the sum is 8. Find the value of a+b when x = 4.
Answer by ankor@dixie-net.com(22740)

A quantity a varies directly as x
a = kx
and another quantity b varies inversely as x.
b = k/x
when x = 2, then sum of a and b is 7,
a + b = 7
Replace a and b
kx + k/x = 7
replace x with 2
2k + k/2 = 7
find k, multiply equation by 2
4k + k = 14
5k = 14
k = 14/5
k = 2.8
:
and when x = 3, the sum is 8.
2.8x + 2.8/x = 8
2.8(3) + 2.8/3 = 9.333 !!!
I just can't make this work

Find the value of a+b when x = 4.